Scandinavian Influence in English? On the Basis of Beowulf and Cursor Mundi

Judit Farkas

28 Judit Farkas

In the history of the English language a widely debated problem is the Scandinavian influence—whether it was as strong and important on the language as it was politically or it had no significant impact on the structure of the language apart from the loanwords the majority of which appeared in geographical and personal names. In fact, as many scholars (cf. e.g. Ekwall (1924, 1963); Wall (1898); Stenton (1948); etc.) have already established, it is extremely difficult to examine and see the impact of the Scandinavian languages on English because of the very similar structure and phonology of these two branches of Germanic. It is especially debated in the Old English period. However, it is a historical fact that there was a strong Viking invasion already in the later 9th century, when the Viking hordes attacked the coasts of Northumbria. The result was the destruction of not only Northumbria but Mercia and East Anglia as well. The Treaty of Wedmor (878), the Danes were compelled to become Christians, was followed by a second treaty, which fixed the Danelaw, i.e. the eastern half of England got under Danish rule. However, these Danes became almost as good Englishmen as the rest. But it took a long time during which the Danes raided England several times, which kept the country in a constant state of alarm. At the end of the 10th, beginning of the 11th centuries this Danish expansion grew in force. The result was that the English were vanquished and Canute became king in England as well as in Denmark (1017–1035). It is true that a new wave of Scandinavian influence reached the western part of England during the 10th and 11th centuries, which is considered to have been quite penetrable and better observable, especially in the vocabulary. We, however, cannot neglect the Danelaw and its predecessors, the Vikings, because in the history of a language if two nations get into direct contact, especially if it is a kind of oppressor–oppressed situation, we can follow the impact of one language on the other, which leads not only to vocabulary changes, which is the easiest to observe as well as to influence, but also to grammatical changes. Better to say, the influence of one language on the other can be measured on the basis of the types of changes it resulted in the languages individually. It is a very well established fact: the bigger is the effect of one language on the other, the more structural, grammatical changes occur in the effected language, because it is grammar and mainly syntax which change the least readily and the most slowly.

It is the aim of the present paper to try to point out the extent of Scandinavian influence in a minor grammatical category, the prepositions with the help of parallelling two extended literary pieces selected from two different historical periods. These periods were different not only from the point of view of the development of the English language, but in the respect of Scandinavian influence as well. The two literary pieces are Beowulf (10th c.) and the Crusor Mundi (14th c.).^{1}

A short matter has to be explained here: why should the prepositions be chosen for such an analysis?

As it was established above, it is grammar and syntax which are the least ready spheres of language to change. We also know that the role of prepositions grew significantly in the course of time, parallel with the simplification of the morphological system of English. Therefore, I thought, it was a category which deserved more interest: if we could point out any Scandinavian influence on the prepositions as early as the 10th century, i.e. a grammatical influence on a minor grammatical category (and especially in a poem which has a clear-cut Scandinavian topic), it would prove our hypothesis that the two languages had been in much closer contact than it has been given attention to so far.

The second part of my comparison is a much later literary piece, the Cursor Mundi from the 14th century. I chose the C version (MS Cotton-Vespasianus, British Museum), which is considered the northernmost version from among the many. The simple reason for this choice was that the time-gap between the two pieces is significant enough—500 years or so—, moreover, this later document comes from a period when there was no direct and strong Scandinavian connection or influence between the two nations. So we may put it “before and after” the bulk of Scandinavian influence in English. Also structurally, it is already a period when the role of prepositions grew considerably. (It is clearly demonstrated by the number of such constructions in the two texts: 588 in Beowulf and 989 in the Cursor, which is almost double so many as in the first text picking out 1650 lines of both. This amount of lines is half of the length of Beowulf (3182), and the same amount is taken from the Crusor Mundi by random choice.)

It might be an interesting study and contrast from another point of view, too. Beowulf was studied in the respect of Scandinavian literary and name elements by many prominent scholars (cf. e.g. Björkman (1910, 1920); Neumann (1912); Sarazzin (1886, 1910); Wessén (1927, 1965); etc.) as the story takes place in Denmark and Southern Sweden. Also the characters are mostly Scandinavians. But the Crusor Mundi has a totally different topic — a highly Biblical one (which assumes Latin impact, if any) — with no Scandinavian elements at all, textually. So it may be quite interesting to see how much a Scandinavian topic was influence by a Scandinavian linguistic impact, and how much it was felt in a totally different topic. Namely, whether the linguistic impact of one language on the other appears already at the time of the strong connection and influence in the written language, or it is still more observable in later times (especially if we do not forget that the Crusor Mundi was written in a period of another significant foreign influence, the Norman French).

Now let us embark on the detailed discussion of what I have found. As it was mentioned above, I have studied 1650 lines of both epic poems and made statistics of the prepositions (see Appendix). Here follows only those details which are of interest from a Scandinavian point of view.

In Beowulf three prepositions deserve interest from the point of view of meaning, because their usage is very much different from the present-day English equivalents. The prepositions are: on, mid and wi�.

On seems to be the most commonly used preposition in Old English having the function of several later prepositions. It was mainly used in the place of present-day English ‘in’. Whenever it stood with the dative case, it most commonly had a meaning of ‘in/at’ with reference to place (55 instances), to time (23), expressing motion ‘in/into’ (5), and a very strong figurative sense usually with reference to locality (27). (This usage can also be parallelled in ON å, på ‘on’, as etymologically both prepositions go back to a common OTeut. ana, an < IE, e.g. å någons wägnar ‘in sy’s name’.)^{2} Once it has the meaning ‘among’, which may have been a further transference of ‘in’. There are only 23 instances where it has the sense of present-day English ‘on’. Whenever it requires the accusative case of the corresponding noun, it has the following senses: ‘into, onto, to’ (52), ‘in’ with time reference (3), ‘away’ in a semi-adverbial phrase (1), and as a required case of the preceding verb (1). With the accusative it usually has some hints at a kind of movement/motion.

Mid is considered to be the original preposition expressing present-day English ‘with’ in Old English. In the majority of cases it stands together with the dative case and it has several meanings: ‘together, along with’ (16), ‘among’ (7), ‘at’ with time reference (1), ‘with’ (12), ‘against’ (1). Whenever it requires the accusative case, it has the sense of ‘with’ (4).

Wi� originally had another meaning than present-day English with. In Beowulf we can find the following senses: with the dative, ‘against’ (5), ‘towards’ (1), ‘from’ (1), ‘with’ (11); with the accusative, ‘against’ (6), ‘towards’ (3), ‘with’ (2).

Etymologically, the picture is the following: Wi�: OE: denoting opposition; denoting accompaniment or addition; de-

noting instrumentality, causation, agency < IE: *wi- denoting separation or division. (Gmc prep. + shortening; ON: vi�r).

Mid: < ComTeut: mit < Gk: μϵτά ‘with’, in the 14th c. superseded by with (ON: me�, me�r, etc. ‘with, along with’).

From the above, we can see that both prepositions underwent some changes in their meanings which have some clear Scandinavian parallels. Can we neglect then the assumption that these Scandinavian meanings might have influenced the native (original) meanings? If it was so, is not it a clear example of a close and significant Scandinavian influence affecting grammatical words? This can be supported by the statistical fact, that these two prepositions occur 41 and 29 times respectively, which makes 11.88% of all the prepositional constructions in the text.

In the Crusor Mundi we could find three prepositions which can clearly be attributed to Scandinavian influence, three words which are suspected to have developed under such influence, especially in their meanings, and there is one preposition which must have had Scandinavian influence in its function. Besides, we may assume that one prepositional construction was reinforced by a parallel Scandinavian pattern.

The three loan prepositions are: Fra/fro, which lives only in some Northern dialects today, and also in Scot-

tish. It goes back to an ON frá ‘from’. It is a ME borrowing into English (OED).

Til(l) ‘to’. This preposition has a local (‘to’) as well as a time (‘until’)

reference. In til(l) ‘into’, which is a compound from native in and Scandinavian til. In

fact, it is a hybrid. Today it is also mainly Scottish and dialectal.

The following three prepositions have Scandinavian impact: With, in this form, is a Common Germanic preposition. However, the

meaning and the change of its meaning hint at some Scandinavian influence. Originally it had the sense ‘in exchange, against, etc.’, i.e. chiefly a completely different sense from its present-day meaning. The original position for expressing the present-day sense ‘togetherness, accompaniment, etc.’ was mid, as we have seen above. Actually this preposition was in force up till the 14th century by which time with pushed it out from this sense. One of the possible causes could be the similarity between mid and the Scandinavian med ‘with’. So we must take it as the result of Scandinavian influence.

Utewit ‘without, outside’. It is again mainly Northern dialectal as well as Scottish. Here our suspicion can be based on the mere fact that ut ‘out’ can be found in both languages with the same meaning, but with in English originally meant ‘against’ (see above). In ON it existed in different meanings, among others ‘at, by, beside’, etc., which hints at the meaning ‘outside’ much more than the original English sense. Moreover, the structure of this compounded preposition resembles the typical ON formation (e.g. �ar vi�), i.e. adverbial + preposition. In English the order was the other way round: preposition + adverbial (e.g. OE wi�utan).

Inwit ‘inside, within’. Its story and development must have been parallel with the previous preposition, or it may be an analogical formation on the previous pattern. Nevertheless, the Scandinavian influence cannot be neglected.

The prepositions which must have developed under Scandinavian influence in their functions are:

At used with verbs, expressing the infinitive marker to. The OED says that it was lost in Southern and Western dialects. Hoad’s etymological dictionary (1987) states that originally this is an Old Teutonic preposition denoting motion and position (‘towards, further’) but this preposition was lost in the above-mentioned dialects and was replaced by to just like in the Scandinavian dialects the to form was lost and the at form preserved. This statement very much supports the Scandinavian influence in the English usage, especially as the infinitive marker in the present-day Scandinavian languages is at/att.

For to + inf. ‘in order to, to’. If we take e.g. Swedish,^{3} the parallels can very well be demonstrated: PrE we went to buy sg, whereas in Swedish we have vi gick för att köpa ngt, which is the equivalent of the for to constructions.

If we make a statistics about the amount of Scandinavian prepositions and influence on prepositions or structures in these randomly chosen 1650 lines, we get the following result: fra/fro: 40 occurrences, til(l): 31, in til(l): 3, with: 86, utewit: 3, inwit: 1, at + inf.: 7, for to + inf.: 32. Total of occurrences: 203, which is 20.52% of the whole amount of prepositions in these lines. It is a considerable amount especially if we keep in mind that we are examining only one particular grammatical means.

Some significant points need to be mentioned, however, to bring this analysis to its conclusion:

(1) The linguistic impact of a language on another can be much better measured in written documents of later periods than in the contemporary ones. It is not at all surprising if we keep in mind that the spoken language adopts neologisms much more quickly and easily than the written language. (Moreover, we must not forget the fact, that these are not holograph pieces, only copies.)

(2) Even though that we could find only two prepositions which have— in my view — clear Scandinavian hints in their semantic changes, I would consider the Scandinavian influence on English as early as the 9th century very significant. If we examine the other prepositions (e.g. in, on, up, etc.) which are the commonest in Beowulf, we can see that they existed in the ON language too, and had similar meanings and functions as in English. If it is so, they must have had their impact on English as far as they strengthened the native usage and they might have led to a much more extended usage as well. Actually, it has been a blank so far. Research neglected the prepositions which are parallels in both languages in contact with the same, or nearly the same meanings. It is an essential fact, however, that such parallels helped and strengthened each other in the different languages. Because in OE and ON there were quite a few of such prepositions, they must be considered too, when we want to measure the impact of one language on the other. Such a consideration will alter the previous results, which were based merely on the borrowed words and foreign meanings. These parallelisms are as important as the studied problems and they should direct attention towards a so far neglected area, namely, the role of parallelisms in measuring mutual impact. On the other hand, it also would underline our initial supposition about a mutual dialect in these English territories rather than different languages. Considering the Crusor Mundi, it appeared in a period when—be-

side the Scandinavian influence—another very significant influence affected the language, the Norman French influence. It had been over by the time when this extended epic was compiled. Thus, we can consider the Cursor Mundi as a kind of synthesis of earlier foreign influences on the language. As the prepositions form a very special group of grammatical words, it must have been the last field which was affected by any foreign influence. In the examined lines and among the 989 prepositional structures we could not point out any one of French origin (neither of Latin, see Appendix), but we found some of possible Scandinavian origin and some which developed under such influence in their meanings or function. Considering this fact, I must conclude this study with my firm assumption that the kind of English–Scandinavian linguistic contact needs much more attention and study, however difficult it might prove to be because of the lack of contemporary written documents as well as of the so widely declared “excuse” that it is impossible because of the similarities of the two languages.

Here emerges a further question—perhaps the topic of another examination: if these two languages, i.e. 9–10th century OE and ON are structurally so close to each other that it is impossible to make a difference between the two considering their mutual impact, can we really speak about:

(a) two different languages—and we are speaking about the Scandinavian occupied British territories: Eastern Scotland, Northumbria, East Anglia—or

(b) did these people develop a mutual dialect involving the characteristics of the two native tongues (and the earlier mentioned parallel prepositions underline this theory), or

(c) their languages, i.e. the dialects of OE spoken in these territories and the dialect of ON spoken by the Scandinavians who settled down here, were in the same kind of relationship as that is among the modern Scandinavian languages, mutual understandability? If it was so, can we neglect so boldly, as many scholars have done so far, this relationship, or do we need to set up a new typology among the Germanic languages and speak rather about different dialects than different languages at this early stage of their development?

34 Judit Farkas

appendix

Beowulf Total number of prepositional constructions: 588

dat  acc  other cases
æfter:  31  inst: 1
æt:  30
be:  10
beforan:  1
buton:  2
for: 24 2
fram/from:  12
in:  26
mid: 37 4
of:  17
on:  124  63
ofer:  5  28
ongean:  2
to: 73 7 gen: 3; gerund: 3
�urh  9
under: 18 12
ut of:  1
wi� 18 11
ymb:  16

Cursor Mundi Total number of prepositional constructions: 989

abote:  3  again(st):  9  amang:  2
apon:  5  apon to:  1  at:  23
be:  6  before:  6  (by)tuix:  12
efter:  12  for:  44  for to:  32
for wit:  3  fro/fra:  40  gain:  1
in:  15  in to:  10  in till:  3
in-wit:  1  o/of:  337  on:  37
ute/out of: 8  overal:  1  ouer/over: 7
til(l):  31  thoro/thoruh: 12  to:  70
under:  4  unto:  5  utewit:  3
wit(h):  86  wit-in/within: 5  wit(o)uten: 28

Scandinavian Influence in English? 35

notes

[1] The sources of the texts I used are Klaeber (1950) for Beowulf and Morris (1962) for Cursor Mundi.

[2] For the Old Norse data, I have been using Gordon (1986).

[3] Cf. Holm & Nylund Lindgren (1977).

references

Björkman, E. (1910) ‘Nordische Personennamen in England in alt- und frühmittelengli-

scher Zeit’. Studien zur Englischen Philologie XXXVII. Björkman, E. (1920) ‘Studien über die Eigennamen in Beowulf’. Studien zur Englischen

Philologie LVIII. Ekwall, E. (1924) ‘The Scandinavian element’. In A. Mawer & F. M. Stenton, edd. Intro-

duction to the Survey of English Placenames. Part I. Cambridge: CUP. 55–92. Ekwall, E. (1963) ‘How long did the Scandinavian language survive in England?’. In:

Selected Papers. Lund Studies in English 33. 65–66. Gordon, E. V. (1986) An Introduction to Old Norse (2nd ed.) Oxford: OUP. Hoad, T. E. (1987) The Concise Oxford Dictionary of English Etymology. Oxford: OUP. Holm, B. and E. Nylund Lindgren (1977) Deskriptive svensk grammatik. Trelleborg. Klaeber, F., ed. (1950) Beowulf and The Fight at Finnsburg. Lexington, MA: D. C. Heath

and Co. Morris, R., ed. (1962) Cursor Mundi. Early English Text Society 99. Oxford: OUP. Neumann, H. (1912) Altnordische Namenstudien. Berlin. OED = The Oxford English Dictionary. Oxford: OUP. 1989. Sarazzin, G. (1886) ‘Altnordisches im Beowulfliede’. Beiträge zur Geschichte der deutschen

Sprache 11. 528–541. Sarazzin, G. (1910) ‘Neue Beowulfstudien IX: Personennamen; Herkunft der Sage.’ En-

glische Studien XLII. 1–37. Stenton, F. M. (1948) ‘The Scandinavian colonies in England and Normandy’. Transac-

tions of the Royal Historical Society 4th Series, vol. 17. London. Wall, A. (1898) ‘A contribution towards the study of the Scandinavian element in the

English dialects’. Anglia 20. 45–136. Wessén, E. (1927) ‘Nordiska Namenstudien’. Uppsala Universitets Årskrift. Wessén, E. (1965) Svensk Språkhistoria III. Stockholm: Alquist & Wiksell.